OpenStudy (anonymous):

What value of x makes the equation true? 6.7 = –3.2 + x A. –9.9 B. –3.5 C. 3.5 D. 9.9

OpenStudy (anonymous):

i think C

OpenStudy (rushwr):

6.7 = -3.2 + x x = 6.7 + 3.2

OpenStudy (rushwr):

therefore x = 9.9

OpenStudy (rushwr):

So D

OpenStudy (anonymous):

i have 3 more

OpenStudy (bishoppatton):

Yes

OpenStudy (rushwr):

open up new questions for those :)